Multiply. (3x – 1)(2x + 5)
Answer by ankor@dixie-net.com(15624)

Multiply. (3x – 1)(2x + 5)
:
Learn the FOIL method which means:
First: 3x * 2x = 6x^2
Outer: 3x * 5 = 15x
Inner: -1 * 2x = -2x
Last: -1 * 5 = -5
:
Put it all together:
:
6x^2 + 15x - 2x - 5
:
NOte the middle terms are like terms, do the math and you have:
6x^2 + 13x - 5:
:
(3x -1)*(2x + 5) = 6x^2 + 13x - 5
:
You must know this method to succeed in algebra.